OpenStudy (anonymous):

whats squre root of 292

OpenStudy (shane_b):

You have access to a computer but not a calculator? :) It's about 17.

OpenStudy (theeric):

@Shane_B , the goal might be to do it without a calculator!

OpenStudy (anonymous):

theric is right

OpenStudy (theeric):

Someone on here suggested you use a factor tree for problems like these. Start with your 292, and divide it by 2, then 3, then 5, then 7 and other primes until one gives you a whole number. Then do the same thing with that number! If you want to be quick, maybe you can look for some different multiples. But when you get to a prime number, you know you're done. 292 2 146 2 73 \(73\) is prime. Now you know that \(2\times 2\times 73 = 292\) Now, what is \(\sqrt{2\times 2\times 73}\)? Well, here's your hint: \(2\times 2 = 2^2\), and \(\sqrt {2^2}=\pm 2\). Also, \(\sqrt{2\times 2\times 73}=\sqrt{2\times 2}\times\sqrt{73}\)

OpenStudy (theeric):

Any questions?

OpenStudy (anonymous):

146

OpenStudy (shane_b):

@theEric: If you saw his other questions you'd see that he doesn't appear to be looking for explanations...just mathematical answers. @danielrrethatoku: Please use the Mathematics group for math questions.

OpenStudy (theeric):

@Shane_B , thank you for the heads up. And yes, @danielrrethatoku , it's easy to accidentally post in the wrong section. But I agree with Shane_B , you should definitely post questions like this in the math section! It's okay to close the question in the wrong section, and post it in the right section. I've seen a moderator encourage that :) I hadn't noticed it was the wrong section! I was jumping around too much. Take care, both of you!